OpenStudy (anonymous):

arcsin __ = arctan (x/x+1)

OpenStudy (anonymous):

x/(2x^2+2x+1)^2

OpenStudy (anonymous):

Put it in that blank

OpenStudy (anonymous):

Sqrt?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

the opposite side is x, adjacent side is x+1 and so, hypotenuse is (2x^2+2x=1)^2

OpenStudy (anonymous):

sqrt(2x^2+2x+1)

OpenStudy (anonymous):

haha its square root not squared

OpenStudy (anonymous):

Oh yes..its square root

OpenStudy (anonymous):

so \[\frac{ x }{ \sqrt{2x ^{2}+2x+1}}\] is the answer?

OpenStudy (anonymous):

Yup @ambernicole u r correct